Anne and Ben have the total weight of 120 kilograms. Ben and Charlie have the total weight of 130 kilograms. Anne and Charlie have the total weight of 150 kilograms. What is the weight of Charlie?
Answer by ankor@dixie-net.com(22740)

Write an equation for each statement:
:
Anne and Ben have the total weight of 120 kilograms.
a + b = 120
b = (120-a)
:
Ben and Charlie have the total weight of 130 kilograms.
b + c = 130
Replace b with 120-a
120-a + c = 130
-a + c = 130 - 120
-a + c = 10
c - 10 = a
:
Anne and Charlie have the total weight of 150 kilograms.
a + c = 150
replace a with (c-10)
(c-10) + c = 150
2c = 150 + 10
2c = 160
c = 160/2
c = 80 kg is Charlie's weight